# Akshay Krishnan1, 2  Xinchen Yan1  Vincent Casser3  Abhijit Kundu1 1Google DeepMind  2Georgia Insitute of Technology  3Waymo

### v1

@@ Abstract
Diffusion models are state-of-the-art for image generation. Trained on large datasets, they capture expressive image priors that have been used for tasks like inpainting, depth, and (surface) normal prediction. However, these models are typically trained for one specific task, e.g., a separate model for each of color, depth, and normal prediction. Such models do not leverage the intrinsic correlation between appearance and geometry, often leading to inconsistent predictions.
In this paper, we propose using a novel image diffusion prior that jointly encodes appearance and geometry. We introduce a diffusion model Orchid, comprising a Variational Autoencoder (VAE) to encode color, depth, and surface normals to a latent space, and a Latent Diffusion Model (LDM) for generating these joint latents. Orchid directly generates photo-realistic color images, relative depth, and surface normals from user-provided text, and can be used to seamlessly create image aligned partial 3D scenes. It can also perform image-conditioned tasks like monocular depth and normal prediction jointly and is competitive in accuracy to state-of-the-art methods designed for those tasks alone. Lastly, our model learns a joint prior that can be used zero-shot as a regularizer for many inverse problems that entangle appearance and geometry. For example, we demonstrate its effectiveness in color-depth-normal inpainting, showcasing its applicability to problems in sparse-view 3D generation. See orchid3d.github.io for more results.
@@ 1 Introduction
Diff + FF Diff + Diff Joint Diff
@@
Inference speed [Uncaptioned image] (mid) [Uncaptioned image] (slow) [Uncaptioned image] (fast)
@@
Table 1: Comparison of methodologies for 2.5D/3D generation: Diff + FF diffuses color image followed by feedforward models for depth + normal prediction; Diff + Diff diffuses images, then diffuses depth and normal conditioned on the color image; Joint Diff (Orchid) jointly diffuses color, depth and normal.
Generative 3D scene modeling [52, 21, 47, 17] has been a longstanding challenge in the computer vision and robotics community for decades. In particular, generating the appearance and geometry of 3D scenes has important applications in virtual reality [46, 51], animation [44], and robotics [14, 1, 5]. Since the scale of available 3D data is small, 3D generation models rely on priors learned from large scale 2D (color) images. In recent years, diffusion models trained on billions of color images have been adopted as an appearance prior for generating 3D objects or scenes. A typical 3D scene generation workflow involves using a standard image diffusion model for appearance followed by separate models each for depth and surface normal estimation (e.g., DepthAnythingV2 [60], Marigold [25], Lotus-G [18], and DSINE [2]). However, the appearance and geometry of 3D scenes are tightly coupled by nature. In this paper, we introduce a joint generative model of appearance and geometry for 3D scene generation.
We propose a new method, Orchid, for the joint diffusion of appearance and geometry in a unified image latent space. Leveraging 2D priors from large-scale pre-training on color images, we propose a new joint latent space that extends the image diffusion latent space to incorporate 3D priors in the form of depth and surface normals. Specifically, we train a Variational Autoencoder (VAE) by taking depth and normal as extra input and output channels, initialized with a pre-trained color space VAE learned from billions of color images. Such training enables us to jointly encode and decode image, depth, and normal using a single model. Given the joint latent space as constructed by our VAE, we further train a latent diffusion model (LDM) to denoise the joint latents on a large dataset. Specifically, we craft a dataset containing coupled color, depth, and normal in a similar fashion as [59, 60] for joint modeling. Our learned joint diffusion priors can be used to directly generate color, depth and surface normals from input text prompts. The same priors can be further used to jointly generate depth and surface normals from input color images by finetuning as a color-conditioned model without text prompts.
A joint prior has several advantages over a combination of an color image prior and color-conditioned downstream models for depth or normal, as summarized in Table 1. It is faster as it only needs a single model. By directly sampling from the joint distribution, it also avoids the compounding error problem of stacking specialized models (performing depth/normal estimation based on generated color output) and the inherent information loss between the models which can cause ambiguity and estimation errors. But most importantly, we found that depth and normal predictions from Orchid are more consistent with each other than depth and normal produced by separate models.
In addition to the above advantages for geometry generation, our joint diffusion model is a 2.5D foundation model that can be used for various applications. We demonstrate that it can be used as an color-conditioned depth and normal prediction model. Incorporating color, depth and normal into a shared latent space also allows us to perform multi-modal inpainting: joint generation of all modalities using a partial subset of observations. Orchid has several potential applications in robotics and virtual reality, including depth completion from color image and sparse LiDAR depth, inpainting holes produced during novel view synthesis from sparse views, and interactive 3D experience. It could also be used in a score-distillation framework, providing a depth and normal aware score function.
In summary, our contributions are as follows:
@@
  • We introduce Orchid, a single joint diffusion model to generate both appearance and geometry. Key to our method is a new joint latent space that extends the image diffusion latent space to incorporate depth and normals.
  • We train Orchid’s VAE with extra input and output channels to learn a joint latent space from a large dataset with paired color, depth, and normals. When trained on the new joint latent space, Orchid can accurately generate appearance and geometry jointly from input text prompts.
  • Orchid serves as a foundation model for various downstream applications, including monocular depth and normal prediction (with finetuning) and zero-shot multi-modal inpainting. Orchid’s monocular depth and normal predictions are competitive with SOTA methods that use separate models, while being more consistent.
@@
Figure 2: Orchid training architecture: As illustrated on the left figure, we first train a joint color-depth-normal VAE that leverages priors from color-only pretraining. We introduce additional depth and normal reconstruction losses, along with a distillation loss to ensure that the joint latent space follows the structure of the original color-only VAE. As illustrated on the right figure, we then train the latent diffusion model on paired image-depth-normal and text data, while keeping the joint VAE encoder frozen.
@@ 2 Related work
Monocular depth estimation is an extensively studied problem in computer vision. Modern deep depth prediction models rely on priors learned from large scale datasets [39, 30, 3, 11, 12, 25, 13, 59, 60, 22]. Since metric depth is a function of the camera’s intrinsics parameters, most approaches predict affine-invariant depth. Models that predict metric depth either predict depth for a single camera [11] or condition on the camera intrinsics [3, 22, 16]. However, since the amount of depth data with known camera intrinsics is limited, most recent methods with zero-shot generalization capabilities, much like our model, predict affine-invariant relative depth.
Monocular surface normal prediction has also seen similar trends as depth prediction, with training deep models for feed-forward prediction being the standard approach [54, 9]. In particular, [9] achieved state-of-the-art performance following this setup. The most recent state-of-the-art [2] incorporates inductive biases specific to the normal prediction task into the deep model.
@@
Most recent approaches for depth [59, 60] and normal prediction [22] adopt self-supervised ViT backbones [37], and train jointly on several large datasets to achieve zero-shot in-the-wild generalization. Although they can predict depth and normals from images, they are discriminative and do not learn a generative prior that can be leveraged for tasks like depth completion.
@@
Several recent works [25, 13, 48, 49, 8, 24] have shown that diffusion priors learned from large color image datasets can be adapted for depth and normal estimation by finetuning them on relatively small amounts of data. These color-conditioned depth and normal diffusion models learn a generative 2D prior of depth which can be used for tasks like depth completion.
However, when used to generate 3D scenes, they impose a two-step process where a color image is first generated using a diffusion model, followed by depth and/or normal generation using another diffusion model. In contrast, our model learns a joint prior over color, depth and normal.
@@
The 2D priors learned by text-to-color diffusion models have been leveraged for controlled generation tasks such as inpainting or super-resolution. The two main approaches to use diffusion priors either 1) train a masked image conditioned model [61] model on additional data [56], or 2) use the conditioning image to guide the diffusion process without additional training [35, 43, 42, 6]. The second approach is more appealing since it does not require additional training and retains the diffusion model’s generalization capabilities. We use this approach with our learned diffusion prior to jointly inpaint color, depth and normals.
The diffusion image priors have also been used for 3D scene generation conditioned on text or single-view color image. One family of approaches uses diffusion models as novel view priors within an optimization framework using a variant of the score-distillation loss [38, 34, 32, 58, 55]. We believe such work can benefit from a joint appearance-geometry prior that avoids the need for cascaded methods and provides score functions that capture the joint appearance-geometry likelihood.
@@ 3 Method
Figure 3: Orchid inference: (a) For text-conditioned generation, We denoise the joint latents conditioned on the text prompt; (b) For monocular depth / normal estimation, we denoise the joint latents with a noise-free latent of the input color image as condition.
Orchid is a latent diffusion model that learns a joint 2D generative prior of appearance (RGB) and 3D geometry (depth and surface normals). More specifically, it learns a score function for the joint distribution $p(\mathbf{x},\mathbf{d},\mathbf{n}|\mathcal{T})$ of color $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, depth $\mathbf{d}\in\mathbb{R}^{H\times W}$, and surface normals $\mathbf{n}\in{[-1,1]}^{H\times W\times 3}$ (with $\mathbf{n}||_{2}=1$), conditioned on text prompts $\mathcal{T}$. This is a foundational generative prior that can be used for various tasks, such as text-conditioned appearance and geometry generation (sampling from $p(\mathbf{x},\mathbf{d},\mathbf{n}|\mathcal{T})$), color-conditioned depth and normal estimation (sampling from $p(\mathbf{d},\mathbf{n}|\mathbf{x})$), and joint multi-modal inpainting.
Central to our method is (1) a joint latent space VAE (Section 3.1) that additionally encodes and decodes depth and normals, building upon 2D priors learned by color-space VAE; and (2) a joint diffusion prior (Section 3.2) with the ability to generate depth and normals in addition to color. The generative priors learn by Orchid can be used for color-conditioned depth and normal prediction by finetuning on an added color image condition (Section 3.3).
@@ 3.1 Joint latent space VAE
The conventional color-space VAE encoder [27, 41] projects pixels to a compact latent space, which diffusion models use for generation by iterative denoising. To jointly model appearance and geometry, we extend the conventional VAE to include depth and normals by adding four more input and output channels (one for depth and three for normals). This design allows Orchid to decode all three modalities jointly at once, without relying on separate models (e.g., monocular depth estimators) to predict geometry. Using a joint VAE for appearance and geometry also allows the encoder to learn a unified compact latent space, minimizing redundancies in depth and normal representations. This is unlike other diffusion-based depth models that duplicate depth across 3 channels to reuse the VAE [25, 13, 18].
We preprocess depth before it is input to the VAE. In particular, from metric depth $\mathbf{d}$, we compute inverse depth $\mathbf{d^{*}}=1/{\mathbf{d}}$, and its mean deviation around the median: $d_{\sigma}=\text{mean}(|\mathbf{d^{*}}-\text{median}(\mathbf{d^{*}})|)$. We normalize inverse depth by the deviation $\mathbf{d^{\prime}}=\mathbf{d^{*}}/d_{\sigma}$. We then shift it to begin at zero: $\mathbf{d}_{\text{model}}=\mathbf{d^{\prime}}-{d^{\prime}}_{\text{min}}$. To simplify notation, we use $\mathbf{d}$ to denote the preprocessed depth $\mathbf{d}_{\text{model}}$.
@@
  $\theta_{\text{enc}}\stackrel{{\scriptstyle+}}{{\leftarrow}}-\nabla_{\theta_{\text{enc}}}(L_{\mathbf{x}}+L_{\mathbf{d}}+L_{\mathbf{n}}+\alpha L_{\text{KL}}+\beta L_{\text{distill}})$
@@ 3.2 Joint diffusion prior for color-depth-normal
We train a latent diffusion model (LDM) to jointly denoise color-depth-normal latents $\mathbf{z}$. In the forward process, we inject noise in the joint latent space through $\mathbf{z}_{t}=\sqrt{\overline{\alpha_{t}}}\mathbf{z}_{0}+\epsilon\sqrt{1-\overline{\alpha_{t}}}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ by following $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{z}_{t};\sqrt{\overline{\alpha_{t}}}\mathbf{z}_{0},(1-\overline{\alpha_{t}})\mathbf{I})$. Here, $\mathbf{z}_{t}$ represents the noisy input at diffusion time $t$ and $\overline{\alpha_{t}}=\prod_{s=1}^{t}(1-\beta_{s})$ is the diffusion coefficient [20]. In the reverse process, our latent diffusion model (parameterized by $\theta_{\text{LDM}}$) predicts the target $y$ given $\mathbf{z}_{t}$ and diffusion time $t$. We optimize the standard objective function $L_{\text{LDM}}=\mathbb{E}_{\epsilon\in\mathcal{N}(0,I),t\in\mathcal{U}(T)}[\|y-f_{\theta_{\text{LDM}}}(\mathbf{z}_{t},t)\|^{2}]$. We use the $v$-prediction (velocity) parametrization for our target $y$ [45]. We train our LDM to denoise joint latents conditioned on text embeddings (Figure 2). This allows us to directly use Orchid to generate color, depth and surface normals from text prompts with classifier-free guidance [19] (Figure 3a).

### v2

# Orchid: Image Latent Diffusion for Joint Appearance and Geometry Generation

@@ Abstract
We introduce Orchid, a unified latent diffusion model that learns a joint appearance-geometry prior to generate color, depth, and surface normal images in a single diffusion process. This unified approach is more efficient and coherent than current pipelines that use separate models for appearance and geometry. Orchid is versatile—it directly generates color, depth, and normal images from text, supports joint monocular depth and normal estimation with color-conditioned finetuning, and seamlessly inpaints large 3D regions by sampling from the joint distribution. It leverages a novel Variational Autoencoder (VAE) that jointly encodes RGB, relative depth, and surface normals into a shared latent space, combined with a latent diffusion model that denoises these latents. Our extensive experiments demonstrate that Orchid delivers competitive performance against SOTA task-specific methods for geometry prediction, even surpassing them in normal-prediction accuracy and depth-normal consistency. It also inpaints color-depth-normal images jointly, with more qualitative realism than existing multi-step methods.
@@
11footnotetext: Work done as an intern at Google DeepMind
@@ 1 Introduction
Imagine that you are planning to paint a beautiful picture of a café. You would not simply start painting—you would first sketch out the space, mapping the layout of seating areas, considering the slopes of the floor and walls, and capturing how light and shadow create the illusion of depth. This natural, integrated understanding of appearance and geometry underpins creative processes in almost every domain from architecture to video game development. Similarly, this joint prior of appearance and geometry is key for deep generative models to produce seamless 3D scenes, opening up transformative possibilities in VR [47, 52], animation [45], and robotics [14, 1, 5]. Although current 2D diffusion models excel at creating high-quality images from a rich appearance prior, we still face challenges in jointly modeling appearance and geometry, largely due to the limited scale of 3D datasets.
Multi-step generation Joint Diff
Diff + FF Diff + Diff (ours)
@@
Inference time (s / image) 1.3* [Uncaptioned image] 4.2* [Uncaptioned image] 1.2 [Uncaptioned image]
@@
Table 1: Comparison of methodologies for 2.5D/3D generation: Diff + FF diffuses color image followed by feedforward ($F$) models for depth + normals; Diff + Diff diffuses color, then diffuses depth and normals conditioned on the color image; Joint Diff (Orchid) jointly diffuses color, depth and normal. Methods with (*) cannot store all models on a GPU and need added I/O time not included here (more details in Appendix Section B).
A viable option for the task of generating images along with its 3D geometry is to first generate a color image using a standalone color image generation model, followed by separate monocular depth and normal models. However, using separate task-specific models for depth and normal results in inconsistencies between their predictions. The overall generation system also becomes inefficient by using multiple large models, with high memory consumption and latency. In addition, this approach of first sampling from an appearance prior, and subsequently sampling from separate color-conditional depth and normal priors does not yield good results for inverse problems that entangle appearance and geometry - such as 3D inpainting, scene completion, or object manipulation. These problems require sampling from the joint space, given some partial observations (for example, a partial colored pointcloud). A method that first samples from a color-image prior using an inpainting model ignores the known scene geometry.
In this work, we propose “Orchid”: a unified latent diffusion model (LDM) for the joint generation of appearance and geometry. Orchid provides a joint generative latent prior that can be readily applied to various problems, avoiding the need for combinations of different task-specific priors. It comprises a new Variational Autoencoder (VAE) to encode RGB, depth, and surface normals to a joint image latent space, and a text-conditioned LDM to denoise the joint latents. This design allows Orchid to address different 3D generation problems. Specifically, it can readily generate color, depth and normal images from text. It is the first model to jointly inpaint large 3D regions in the color-depth-normal image space. With color-conditioned finetuning, it can jointly predict depth and normals from input color images, acting as a monocular geometry estimator. Both our VAE and our LDM leverage pretraining from large color image datasets. Additionally, we use multiple real world color-depth-normal datasets, along with distillation from discriminative teacher depth and normal models on a large text-image dataset.
As shown in Table 1, a joint LDM has several advantages over a combination of a color LDM and color-conditioned downstream models for depth or normals. It generates color, depth, and normals in a single diffusion process, being faster and more memory efficient because only a single model is needed. As shown by our experiments, its depth and surface normal predictions are more self-consistent than those produced by different state-of-the-art models. Finally, the joint prior can be directly applied to many 3D reconstruction problems: we show that it can be used to sample from the joint appearance-geometry distribution conditioned on partial input, such as during 3D scene inpainting. Orchid has several potential applications in robotics and virtual reality, such as image-conditioned depth completion for sparse LiDAR sensors or inpainting unseen regions during novel view synthesis from sparse views. It could also serve as a replacement for separate appearance-geometry models used in score-distillation frameworks such as [65] by providing a unified appearance-geometry score function.
In summary, our contributions are as follows.
@@
  • We introduce Orchid, a unified latent diffusion model to generate color, depth and normals from text in a single diffusion process. We train a new joint VAE to encode color, depth, and normals to a shared latent space.
  • We show that Orchid can act as a joint monocular depth and normal estimator with color-conditioned finetuning. Orchid beats state-of-the-art methods in terms of monocular normal prediction accuracy and consistency of depth and normal predictions.
  • Orchid learns a joint color-depth-normal prior that can be used for inverse problems. We show that is can consistently inpaint 3D scenes in the color-depth-normal space, producing qualitatively better results than a combination of color, depth and normal inpainting models.
@@
Figure 2: Orchid training scheme: As illustrated on the left figure, we first train a joint color-depth-normal VAE that leverages priors from color-only pretraining. We introduce additional depth and normal reconstruction losses, along with a distillation loss to ensure that the joint latent space follows the structure of the original color-only VAE. As illustrated on the right figure, we then train the latent diffusion model on paired image-depth-normal and text data, while keeping the joint VAE encoder frozen.
@@ 2 Related work
Modern deep depth prediction models rely on priors learned from large scale datasets [41, 29, 3, 11, 12, 24, 13, 63, 64, 21] to generalize beyond training distribution. Since metric depth is a function of the camera’s intrinsics parameters, most approaches, including ours, predict affine-invariant depth. Models do that predict metric depth are either limited to a single camera [11] or condition on the camera intrinsics [3, 21, 17].
Monocular surface normal prediction has also seen similar trends, with feed-forward deep models being the standard approach [57, 9]. In particular, [9] achieved state-of-the-art performance following this setup. The most recent state-of-the-art [2] incorporates additional inductive biases specific to the normal prediction task into the model.
@@
Most recent approaches for depth [63, 64] and normal prediction [21] adopt self-supervised ViT backbones [37], and train jointly on several large datasets to achieve zero-shot in-the-wild generalization. Although they make predictions from images, they are discriminative and do not learn a generative prior that can be leveraged for tasks like depth completion or refinement.
@@
Several recent works [24, 13, 48, 49, 8, 23, 18, 7, 68] have shown that diffusion priors learned from large color image datasets can be adapted for depth and normal estimation by finetuning them on relatively small amounts of data. These color-conditioned diffusion models learn a generative 2D prior which can be used for problems like depth completion. GeoWizard [13] is an existing diffusion model for joint depth-normal prediction, but it cannot generate color. It also uses separate representations with “switch conditioning” for depth and normals, consuming more memory.
When used to generate 3D scenes, these methods require multiple diffusion processes and models: first to generate color, then to generate depth and/or normals. In contrast, our model learns a joint prior over color, depth and normal.
@@
2D text-to-color diffusion models trained on large datasets are emerging as promising generative priors for several applications. They are used as denoisers within an optimization framework, with other auxiliary objectives. They are leveraged for text-to-3D generation using a score-distillation sampling (SDS) loss to optimize a neural 3D representation [39, 33, 31, 62, 58]. They can also be used for 3D object manipulation [56, 60], depth completion [54], and 3D scene completion [59, 51, 40, 35, 65]. The biggest challenge for such methods is the multi-view inconsistency of color diffusion models, as inconsistent appearance predictions across views cause artifacts in 3D reconstructions. These methods only rely on appearance priors. Some works [65] combine color with geometric diffusion priors by balancing multiple diffusion models. Orchid introduces a stronger unified appearance-geometry prior in a single model.
@@ 3 Method
Figure 3: Orchid inference: (a) For text-conditioned generation, we denoise the joint latents conditioned on the text prompt; (b) For monocular depth / normal estimation, we denoise the joint latents with a noise-free input color-only latent as condition (no text).
We train Orchid, an LDM to learn a joint 2D generative prior of appearance (RGB) and geometry (depth and surface normals). Specifically, it learns a score function for the joint distribution $p(\mathbf{x},\mathbf{d},\mathbf{n}|\mathcal{T})$ of color $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, depth $\mathbf{d}\in\mathbb{R}^{H\times W}$, and surface normals $\mathbf{n}\in{[-1,1]}^{H\times W\times 3}$ (with $\mathbf{n}||_{2}=1$), conditioned on text prompts $\mathcal{T}$. This is a foundational generative prior that can be used for various tasks: text-conditioned appearance and geometry generation (sampling from $p(\mathbf{x},\mathbf{d},\mathbf{n}|\mathcal{T})$), color-conditioned depth and normal estimation (sampling from $p(\mathbf{d},\mathbf{n}|\mathbf{x})$), and joint appearance-geometry inpainting (sampling from $p(\mathbf{x},\mathbf{d},\mathbf{n}|\mathcal{T};\mathbf{x^{\prime}},\mathbf{d^{\prime}},\mathbf{n^{\prime}})$ where $\mathbf{x^{\prime}},\mathbf{d^{\prime}},\mathbf{n^{\prime}}$ are partial observations). The ability to condition on partial observations in the joint space is unique to Orchid.
Central to Orchid is (1) a VAE (Section 3.1) that encodes color, depth and normals to a joint latent space, building upon priors learned by color-space VAE; and (2) a text-conditioned LDM (Section 3.2) with the ability to generate depth and normals in addition to color. We also finetune Orchid for monocular depth and normal prediction, with an added color image condition (Section 3.3).
@@ 3.1 Joint latent space VAE
The VAE encoder [26, 44] used by color LDMs [43] projects RGB images to a compact latent space for iterative denoising. We extend this VAE by adding four channels (one for depth and three for normals) to jointly encode color, depth and normals. This lets Orchid decode all three modalities at once, unlike other task-specific LDMs [24, 13, 18] that use separate latents and LDMs for each modality. The joint latent space also minimizes redundancies in depth and normal representations. Moreover, training our VAE enables us to use pixel space representations that are best suited for geometry. For e.g., using unbounded inverse depth, which yields better predictions at longer ranges, unlike models that reuse a color-VAE with a [0-1] depth normalization using near-far planes.
We preprocess depth before it is input to the VAE. In particular, from metric depth $\mathbf{d}$, we compute inverse depth $\mathbf{d^{*}}=1/{\mathbf{d}}$, and its deviation around the median: $d_{\sigma}=\text{mean}(|\mathbf{d^{*}}-\text{median}(\mathbf{d^{*}})|)$. We normalize inverse depth by the deviation $\mathbf{d^{\prime}}=\mathbf{d^{*}}/d_{\sigma}$. We then shift it to begin at zero: $\mathbf{d}_{\text{model}}=\mathbf{d^{\prime}}-{d^{\prime}}_{\text{min}}$. To simplify notation, we use $\mathbf{d}$ to denote the preprocessed inverse depth $\mathbf{d}_{\text{model}}$.
@@
  $\mathbf{z_{\mu}^{*}},\mathbf{z_{\sigma}^{*}}]\leftarrow\text{Enc}(\mathbf{x},\mathbf{d},\mathbf{n};\theta_{\text{enc}}^{*})$
[…]
@@
  $\theta_{\text{enc}}\stackrel{{\scriptstyle+}}{{\leftarrow}}-\nabla_{\theta_{\text{enc}}}(L_{\mathbf{x}}+L_{\mathbf{d}}+L_{\mathbf{n}}+L_{\text{KL}}+L_{\text{distill}})$
@@ 3.2 Joint diffusion prior for color-depth-normal
We train a latent diffusion model (LDM) to jointly denoise color-depth-normal latents $\mathbf{z}$, conditioned on text embeddings (Figure 2). In the forward process, we inject noise in the joint latent space through $\mathbf{z}_{t}=\sqrt{\overline{\alpha_{t}}}\mathbf{z}_{0}+\epsilon\sqrt{1-\overline{\alpha_{t}}}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ by following $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\mathbf{z}_{t};\sqrt{\overline{\alpha_{t}}}\mathbf{z}_{0},(1-\overline{\alpha_{t}})\mathbf{I})$. Here, $\mathbf{z}_{t}$ represents the noisy input at diffusion time $t$ and $\overline{\alpha_{t}}=\prod_{s=1}^{t}(1-\beta_{s})$ is the diffusion coefficient [20]. In the reverse process, our latent diffusion model (parameterized by $\theta_{\text{LDM}}$) predicts the target $y$ given $\mathbf{z}_{t}$ and diffusion time $t$. We optimize the standard objective function $L_{\text{LDM}}=\mathbb{E}_{\epsilon\in\mathcal{N}(0,I),t\in\mathcal{U}(T)}[\|y-f_{\theta_{\text{LDM}}}(\mathbf{z}_{t},t)\|^{2}]$. We use the $v$-prediction (velocity) parametrization for our target $y$ [46]. This formulation allows us to directly use Orchid to generate color, depth, and normals from text prompts with classifier-free guidance [19] (Figure 3a).